what will be the ratio of mu between two goods at MRSxy=3:1


1
Expert's answer
2021-09-28T16:57:27-0400

"MRS_{xy} = \frac{dy}{dx} = \frac{MUy}{MUx}"

= "\frac{1}{3}"

Therefore, the ratio "MUy:MUx = 1:3"